# Multi-Level Explanations for Generative Language Models

### v1

@@ Abstract
Perturbation-based explanation methods such as LIME and SHAP are commonly applied to text classification. This work focuses on their extension to generative language models. To address the challenges of text as output and long text inputs, we propose a general framework called MExGen that can be instantiated with different attribution algorithms. To handle text output, we introduce the notion of scalarizers for mapping text to real numbers and investigate multiple possibilities. To handle long inputs, we take a multi-level approach, proceeding from coarser levels of granularity to finer ones, and focus on algorithms with linear scaling in model queries. We conduct a systematic evaluation, both automated and human, of perturbation-based attribution methods for summarization and context-grounded question answering. The results show that our framework can provide more locally faithful explanations of generated outputs.
@@ 1 Introduction
Perturbation-based explanation methods, such as LIME [22] and SHAP [14], are popular for explaining individual predictions of black-box classification and regression models. The explanations take the form of attribution scores assigned to each part of the input, quantifying the effect of that part on the prediction when perturbed. [^1] We use the broader term “input attribution” in this paper rather than “feature attribution” because “features” may no longer be the best term for input parts such as text spans. For text input, the input parts are spans of text (often words), and applications to text classification tasks such as sentiment analysis have been well-studied [2, 13, 17, 12].
This work focuses on perturbation-based explanations for generative language tasks such as summarization and question answering. There has been little work on such explanations for text generation as corroborated by a survey paper on SHAP-based methods [17], the most relevant work being a collection of text examples in the SHAP package documentation [25] and not a paper. Concurrent with this work, the Captum library has added new capabilities for generative language models [16].
@@
Perturbation-based input attribution for generative tasks presents challenges related to having text as output and long texts as input. We propose a general framework called MExGen (Multi-level Explanations for Generative Language Models) to address these challenges. The framework can be instantiated with different attribution algorithms, and we do so using SHAP- and LIME-like algorithms.
Figure 1: Example of using MExGen C-LIME (BERT scalarizer) to generate explanations. The example shows how the algorithm captures copied text (further skeletal remains) but also abstract concept mappings (M54 $\rightarrow$ motorway).
@@
Challenge of output text
The first challenge is that language models output text rather than a real number (e.g., the predicted log probability of a class). Attribution algorithms require a real-valued function to quantify that function’s sensitivity to different inputs. We address this through the concept of scalarizers, functions that map output text to real numbers. We investigate the use of multiple scalarizers and compare the resulting attributions.
Importantly, most scalarizers that we consider address the truly “black-box” setting in which we receive only text as output from the model. This is common with large language models (LLM) that only provide API access or are proprietary. In this case, scalarizers can use only the output text from the model being explained. The text-only setting is not handled by Captum and is only briefly alluded to in SHAP [25].
Challenge of input length
The second challenge is that the input text for generative tasks can be long, particularly for summarization and context-grounded question answering. Longer inputs require more resources, in terms of input perturbations and model queries, and impose financial costs for LLMs with API-only access. Long inputs also pose interpretation issues, since it is not clear what level of granularity in the attributions is most useful to a user. We address the challenge in three ways:
Linguistic segmentation. We segment the input text into linguistic units at multiple levels, for example sentences, phrases, and words, taking advantage of linguistic structure present in the text.
Multi-level attributions. We use a refinement strategy that proceeds from attributing at a coarser level like sentences to a finer level like phrases or words, with optional user input. This controls the number and granularity of units, and limits the compute.
Linear-complexity algorithms. We instantiate our framework with attribution methods whose number of perturbations and model queries scale linearly with the number of units at any level (e.g., number of sentences). In the SHAP category, we choose Local Shapley (L-Shapley) [4]. For LIME-type methods, we propose a linear-complexity variant that also limits the number of units perturbed at one time.
We evaluate the instantiations of MExGen on the tasks of summarization and context-based question answering. The automated evaluation shows that instantiations of MExGen are more locally faithful; particularly, they assign the highest importance to the input parts that are more relevant to the model output with higher frequency than the baselines. Automated evaluations also show that different scalarizers may yield similar explanations depending on the model being explained and the task being performed. Specifically, we found that some scalarizers — which only use API access to the model — generate explanations that closely resemble explanations using the log probabilities of the model being explained. Human evaluation results not only corroborated the automated evaluation findings but also provided additional insights into how certain scalarizers or attribution methods, previously considered similar to other methods in automatic evaluation, were perceived as more faithful by users.
Below is a summary of our main contributions:
  • We investigate several scalarizers for mapping output text to real numbers, notably handling the case of text-only output from the model.
  • We conduct a systematic evaluation, both automated and human, of input attribution methods for summarization and question answering, showing that MExGen can provide more locally faithful explanations of generated output than the alternative methods available.
@@ 2 Related Work
Post Hoc Explanations for LLMs
We focus on perturbation-based post hoc explanation methods for LLMs, notably those with only API access where text is the only available output.
Gradient-based methods provide input attribution explanations [27, 29, 26, 1], but they require access to model gradients with respect to the input.
While there is a significant literature on explaining text classification models, the literature on perturbation-based explanations for natural language generation is scarce, as noted in a survey of SHAP-based methods [17]. Below we discuss the two most relevant works that we have found. One common trait is that they require model logits (possibly from a proxy model) to generate explanations.
PartitionSHAP, the default algorithm for text in the SHAP library, handles long inputs by partitioning them into clusters of tokens and attributing the same score to tokens within a cluster. We have not found a paper describing PartitionSHAP. The description of the SHAP library’s text examples [25] is also cursory. However, it is apparent that PartitionSHAP produces one set of attributions for each token in the output. This approach is less interpretable because it requires the selection of an output token and assigns multiple attribution scores to each input span. PartitionSHAP can work with API-only access [24], but understanding how it does so required significant investigation on our part (see Appendix A.1).
CaptumLIME is our name for a modification of LIME [22] tailored for text generation tasks using new features in the Captum library [16]. [^2] Captum also has variants of SHAP but we found them slow to run and obviated by the availability of PartitionSHAP. CaptumLIME allows the user to manually define units for attribution within the input. It handles sequence outputs by computing log probabilities for tokens in a given output and then summing. This does, however, depend on access to output probabilities, so CaptumLIME is not suitable for the text-only setting.
Method LT Output LT Input API Access
LIME
SHAP
HEDGE
P-SHAP
Captum
MExGen
Table 1: Comparing the features (Long Text Output/Input, API access) of our MExGen framework with existing methods.
Hierarchical Explanations
A line of work [28, 11, 2, 12] has developed hierarchical explanations for sequence models, including LLMs, which can reveal compositional interactions between words and phrases. In particular, the HEDGE algorithm of Chen et al. [2] was identified by Mosca et al. [17] as “arguably the most suitable choice” for NLP input attribution, in part because it builds its hierarchy in a top-down, divisive fashion (as opposed to bottom-up agglomeration [28, 12]), which is more practical for long texts. However, HEDGE is specific to classification because it measures feature importance based on classification margin. Table 1 shows that PartitionSHAP, Captum, and HEDGE each lack capabilities compared to MExGen.
Natural Language Self-Explanations
We also note works have utilized the generative model itself to provide explanations in line with subsequent outputs, referred to as chain-of-thought or CoT [33]. These methods have been shown to be both unstable [32] and highly variable in quality [18]. Further, CoT provides self-explanations in natural language, and not input attribution explanations.
@@ 3 Multi-Level Explanations for Generative Language Models
We describe the proposed MExGen framework for input attribution for generative language models. Figure 2 provides an overview.
In the setting of perturbation-based input attribution, we are given a generative LLM $f$, an input text sequence of interest $x^{o}$ (left side of Figure 2, superscript $o$ for “original”), and a generated output $y^{o}=f(x^{o})$ that is also a text sequence (top middle of Figure 2). Our goal is to explain the output $y^{o}$ by attributing to parts of the input $x^{o}$. Each part of the input, denoted $x_{s}$, $s=1,\dots,d$, is to be assigned an attribution score $\xi_{s}$ (represented by color on the right of Figure 2) quantifying the importance of $x_{s}$ in generating the output, in the sense that if important parts are perturbed, then the output will change significantly. As the second through fourth paths in Figure 2 indicate, model $f$ can be queried on perturbations $x$ of $x^{o}$, with no further access.
@@ 3.1 Handling Text Outputs
Input attribution methods require a numerical target function as the object to explain. Since an LLM $f$ may only output text, we introduce scalarizers, which are functions $S$ that map output text back to real numbers (shown as blue boxes in Figure 2). We consider two types of access to $f$: (a) full logit access, where $f$ can provide predicted logits for all tokens in its vocabulary, at each position in the output sequence, (b) text-only setting, where we are limited to text outputs. See Appendix A.1 for why vocabulary-wide access is assumed for (a).
@@
When all logits are available, we use the probability of generating the original output sequence $y^{o}$ as the function to explain. We refer to this as the Log Prob scalarizer. Given output sequence $y^{o}$ of length $\ell$ and an arbitrary input sequence $x$, we compute from the model’s logits the log probability of generating each original output token $y^{o}_{t}$ conditioned on $x$ and previous output tokens $y^{o}_{<t}$. We then average over the output sequence to obtain the scalarized output for $x$,
@@
Using the output sequence $y=f(x)$ generated from the input $x$, we consider similarity measures $S(y;y^{o})$ between $y$ and the original output $y^{o}$ as scalarizers. These now depend on $f$ only via composition, i.e., $S(f(x);y^{o})$, and are not parameterized by $f$ as in (1).
  • Sim: Computes the cosine similarity between embedding vectors for $y$ and $y^{o}$.
  • BERT: Computes the BERTScore [35] between $y$ and $y^{o}$.

### v2

@@ Abstract
Despite the increasing use of large language models (LLMs) for context-grounded tasks like summarization and question-answering, understanding what makes an LLM produce a certain response is challenging. We propose Multi-Level Explanations for Generative Language Models (MExGen), a technique to provide explanations for context-grounded text generation. MExGen assigns scores to parts of the context to quantify their influence on the model’s output. It extends attribution methods like LIME and SHAP to LLMs used in context-grounded tasks where (1) inference cost is high, (2) input text is long, and (3) the output is text. We conduct a systematic evaluation, both automated and human, of perturbation-based attribution methods for summarization and question answering. The results show that our framework can provide more faithful explanations of generated output than available alternatives, including LLM self-explanations. We open-source code for MExGen as part of the ICX360 toolkit: https://github.com/IBM/ICX360.
Multi-Level Explanations for Generative Language Models
Lucas Monteiro Paesthanks: Denotes equal contribution in alphabetical order.1, Dennis Wei11footnotemark: 12, Hyo Jin Do2, Hendrik Strobelt2, Ronny Luss2, Amit Dhurandhar2, Manish Nagireddy2, Karthikeyan Natesan Ramamurthy2, Prasanna Sattigeri2, Werner Geyer2, Soumya Ghoshthanks: Work done while at IBM Research.3 1Harvard University           2IBM Research           3Merck Research Labs lucaspaes@g.harvard.edu, dwei@us.ibm.com
@@ 1 Introduction
Large language models (LLMs) are being deployed to generate text used for decision-making, e.g., summarizing meetings Laskar et al. (2023), extracting key points from legal documents Kanapala et al. (2017), and answering doctors’ questions Xiong et al. (2024). In these applications, the LLM is grounded in context (e.g., legal documents) provided as part of the input. Given an LLM response, users may wish to know which parts of the context were responsible for the response, or whether the response is grounded at all Huang et al. (2023a). We propose Multi-level Explanations for Generative Language Models (MExGen) to fulfill this user necessity by providing explanations for context-grounded text generation.
MExGen computes explanations by comparing multiple model predictions given perturbed versions of the input context. The explanations take the form of attribution scores assigned to parts of the input, quantifying the effect of each part on the model output. Accordingly, we also refer to this type of explanation as input attribution. MExGen generalizes popular perturbation-based explanation methods such as LIME Ribeiro et al. (2016) and SHAP Lundberg and Lee (2017) to generative LLMs. Such methods are widely used for text classification Chen et al. (2020); Kim et al. (2020); Mosca et al. (2022); Ju et al. (2023), but their application to generative LLMs is still limited (Section 2).
@@
Perturbation-based input attribution for generative LLMs presents challenges related to having (1) text outputs, (2) high inference cost, and (3) long text inputs. MExGen provides a framework to address these challenges. The framework can be instantiated with different attribution algorithms, and we do so using LIME- and SHAP-like algorithms.
Figure 1: Explanation generated using MExGen C-LIME (with BERT scalarizer) for a summarization example. The most important phrases found by MExGen (darker blue) suggest that the LLM closely paraphrases text (further skeletal remains) and also abstracts concepts (M54 $\rightarrow$ motorway).
Challenge of output text
The first challenge is that LLMs output text rather than a real number (e.g., the predicted log probability of a class). Attribution algorithms require a real-valued function to quantify that function’s sensitivity to different inputs. We address this through the concept of scalarizers, functions that map output text to real numbers. We investigate and compare multiple scalarizers to provide guidance on this choice.
Importantly, most scalarizers that we consider address the truly “black-box” setting in which we receive only text as output from the model. This is common with LLMs that only provide API access or are proprietary.
Challenge of input length and LLM inference
The input text can be long for context-grounded generative tasks, e.g., whole papers or news articles. Generating explanations for long inputs requires more model inferences, imposing higher computational and financial costs. Long inputs also pose interpretation issues, since attributions that are too fine-grained may be less interpretable to a user.
We address this challenge in three ways: (1) Linguistic segmentation: We segment the input text into linguistic units at multiple levels, for example sentences, phrases, and words. (2) Multi-level attribution: We use a refinement strategy that proceeds from attributing at a coarser level like paragraphs to a finer level like words, only refining the most important parts of the context. This controls the number of model inferences and resembles a binary search. (3) Linear-complexity algorithms: We instantiate our framework with attribution algorithms whose numbers of model inferences scale linearly with the number of units (e.g., number of sentences). We also propose a linear-complexity variant of LIME called C-LIME that limits the number of units perturbed at one time.
Evaluation
We evaluate instantiations of MExGen on summarization and context-grounded QA tasks. We show that MExGen provides more faithful explanations compared to baselines (including LLM self-explanation), assigning higher importance to input parts that have the greatest effect on the model output. We also find that certain scalarizers yield similar explanations to each other. Some scalarizers that only use text generate explanations close to those that depend on log probabilities. Human evaluation corroborates the automated evaluation findings and also indicates that certain scalarizers and attribution methods, previously considered similar to others in the automated evaluation, were perceived as more faithful by users.
Our main contributions are:
  • We compare several scalarizers for mapping output text to real numbers, notably handling the case of text-only output from the model.
  • We conduct a systematic evaluation, both automated and human, of input attribution methods for summarization and QA, showing that MExGen can provide more faithful explanations of generated output than the available alternatives. This advantage extends to self-explanations of LLMs (Huang et al., 2023b; Madsen et al., 2024), even from powerful LLMs such as DeepSeek-V3.
@@ 2 Related Work
We discuss works on perturbation-based explanations and self-explanations for generative LMs [^1] Please see Appendix A for other forms of explanation.. The literature on the former for generative LMs is limited, as corroborated by Mosca et al. (2022).
Perturbation-based methods
PartitionSHAP in the SHAP library handles long inputs by dividing them into token spans and assigning the same score to each token in a span. PartitionSHAP produces separate attributions for each output token. This approach is less interpretable because it requires the selection of an output token and assigns multiple attribution scores to each input span. Although PartitionSHAP supports API-only access SHAP (2024b), understanding its operation required substantial effort (see Appendix B.1).
CaptumLIME is a modification of LIME Ribeiro et al. (2016) tailored for text generation tasks in the Captum library Miglani et al. (2023). [^2] Captum also has variants of SHAP but we found them slow to run and obviated by PartitionSHAP. CaptumLIME allows the user to manually define units for attribution within the input. It handles text outputs by computing the log probability of the output. This does, however, require access to output probabilities, so CaptumLIME is not suitable for the text-only setting.
TextGenSHAP Enouen et al. (2023) offers a more efficient sampling strategy for Shapley value estimation based on speculative decoding Leviathan et al. (2023). As with CaptumLIME, such computations also require access to output probabilities. Also, TextGenSHAP is tailored to SHAP.
Table 3 shows that the above methods each lack capabilities that are offered by MExGen.
Self-explanation methods
These methods prompt the model to explain its predictions (e.g., by ranking the most important parts of its context) Camburu et al. (2018); Huang et al. (2023b); Kroeger et al. (2024); Madsen et al. (2024). However, prior work has found that these self-explanations may be less faithful when used for in-context classification tasks Madsen et al. (2024); Huang et al. (2023b). Fragkathoulas and Chlapanis (2024) demonstrated that self-explanations are not as faithful as numerical attributions (like MExGen) in identifying keywords for a QA task. We generalize their work to additional generative tasks and to self-explanations in the form of rankings.
@@ 3 Multi-Level Explanations for Generative Language Models
This section describes the proposed MExGen framework. Figure 2 provides an overview of MExGen.
In the setting of perturbation-based input attribution, we are given a generative LM $f$, an input text sequence of interest $x^{o}$ (left side of Figure 2, superscript $o$ for “original”), and a generated output $y^{o}=f(x^{o})$ that is also a text sequence (top center of Figure 2). Our goal is to explain $y^{o}$ (the target output for explanation) by attributing to parts of the input $x^{o}$. Each part of the input, denoted $x_{s}$, $s=1,\dots,d$, is to be assigned an attribution score $\xi_{s}$ (represented by color on the right of Figure 2) quantifying the importance of $x_{s}$ in generating the output, in the sense that if important parts are perturbed, then the output will change significantly. As the second through fourth paths in Figure 2 indicate, model $f$ can be queried on perturbations $x$ of $x^{o}$, with no further access to $f$.
@@ 3.1 Handling Text Outputs
Input attribution algorithms such as SHAP and LIME require a real-valued function as the object to explain (see Section 3.2, “Linear-complexity algorithms” for examples of how this function is used). Since the LM $f$ may only output text, we introduce scalarizers, which are functions $S$ (shown as blue boxes in Figure 2) that map output text back to real numbers. We consider two types of access to $f$: (a) full logit access, where $f$ can provide predicted logits for all tokens in its vocabulary, at each position in the output sequence; (b) text-only access, where we are limited to text outputs. See Appendix B.1 for why we assume vocabulary-wide logits for (a).
@@
When all logits are available, we use the probability of generating the target output $y^{o}$ as the function to explain. We refer to this as the Log Prob scalarizer. Given output sequence $y^{o}$ of length $\ell$ and an arbitrary input sequence $x$, we compute the model’s log probability of generating each target output token $y^{o}_{t}$ conditioned on $x$ and previous output tokens $y^{o}_{<t}$. We then average over output tokens to obtain the scalarized output for $x$,
@@
Given only the output text $y=f(x)$ generated from the perturbed input $x$, we consider similarity measures $S(y;y^{o})$ between $y$ and the target output $y^{o}$ as scalarizers. These now depend on $f$ only via composition, i.e., $S(f(x);y^{o})$, and are not parameterized by $f$ as in (1). Appendix B.1 has further details on these scalarizers.
  • Sim: Cosine similarity between embeddings of $y$ and $y^{o}$ (e.g., SentenceTransformers embeddings).
  • BERT: BERTScore Zhang et al. (2020) between $y$ and $y^{o}$.
  • BART: BARTScore (“faithfulness” version) between $y$ and $y^{o}$ Yuan et al. (2021). Measures the probability of an auxiliary LM $f_{\textit{BART}}$ generating $y^{o}$ given $y$ as input.
  • SUMM: Similar to the BART scalarizer with a summarization model as $f_{\textit{BART}}$ SHAP (2024b).
  • Log NLI: Log-odds of entailment given $y^{o}$ as premise and $y$ as hypothesis, computed using a natural language inference (NLI) model.
3.2 Handling Long Text Inputs
Some generative language tasks require long input texts, such as in summarization and context-grounded QA. We address this challenge through a combination of three techniques: segmenting the input into linguistic units at multiple levels, using attribution algorithms with linear complexity in the number of input units, and obtaining attributions in a coarse-to-fine manner.
Linguistic segmentation
We segment the input into linguistic units at multiple granularities: paragraphs, sentences, phrases, and words (“Select Units” box in Figure 2). This approach takes advantage of linguistic and other structure present in the input. For example, the input may already be broken into paragraphs or contain multiple distinct retrieved documents, in which case these paragraphs or documents can form the units at the highest level. In contrast, existing methods rely on the model’s tokenizer, which can yield units (tokens) that are too fine, or treat the text as a flat sequence of tokens and let the algorithm decide how to segment it Chen et al. (2020); SHAP (2024a).
We use tokenization and dependency parsing from spaCy v3.6 Honnibal et al. (2020) to segment paragraphs into sentences and words. To segment sentences into phrases, we implemented an algorithm that uses the dependency parse tree from spaCy. In the first pass, the algorithm recursively segments the tree and its subtrees into phrases that are no longer than a maximum phrase length. In the second pass, some short phrases are re-merged. More details are in Appendix B.2.
Our framework also allows for units at any level to be marked as not of interest for attribution. Textual elements like punctuation, prompt templates, and system prompts are usually not relevant for attribution and, therefore, marked as not of interest. Ultimately, the user may choose which parts of the input context are not of interest. Appendix C.3 gives more details on the choice of ignored units.
Linear-complexity algorithms
Given a segmented input with possibly mixed units as in Figure 2, we have the task of attributing to units $x_{1},\dots,x_{d}$ (the “Post Hoc Explainer” block in Figure 2). Here, we consider only perturbation-based attribution algorithms that scale linearly with the number of units $d$ in terms of model queries, to control this cost. We instantiate MExGen with three such methods: leave-one-out (LOO), a LIME-like algorithm with further constraints (C-LIME), and Local Shapley (L-SHAP). In this work, perturbing a unit means simply dropping it; see Appendix F for further discussion.
LOO: Units $x_{1},\dots,x_{d}$ are perturbed one at a time, yielding corresponding perturbed inputs $x^{(1)},\dots,x^{(d)}$. The attribution score for $x_{s}$ is the decrease in scalarizer score due to leaving $x_{s}$ out: $\xi_{s}=S(x^{o};y^{o},f)-S(x^{(s)};y^{o},f)$.
C-LIME: We use a linear model that operates on interpretable features $z$ and approximates the model $f$ in the vicinity of original input $x^{o}$. In our case, the interpretable features $z\in\{0,1\}^{d}$ correspond to units $x_{1},\dots,x_{d}$, with $z_{s}=0$ if unit $s$ is perturbed and $z_{s}=1$ otherwise. The linear model is fit using $n$ perturbations $x^{(1)},\dots,x^{(n)}$ of $x^{o}$, with corresponding interpretable representations $z^{(1)},\dots,z^{(n)}$, and scalarized model outputs $S(x^{(1)};y^{o},f),\dots,S(x^{(n)};y^{o},f)$:
$\xi=\operatorname*{arg\,min}_{w}\sum_{i=1}^{n}\pi\bigl{(}z^{(i)}\bigr{)}\bigl{(}w^{T}z^{(i)}-S(x^{(i)};y^{o},f)\bigl{)}^{2}\\ +\lambda R(w),$ (2)
We make two main departures from LIME, in addition to the use of a general scalarizer $S$ as seen in (2). Firstly, we limit the number of perturbations $n$ to a multiple of the number of units $d$. Since $n$ is the number of samples in the linear regression problem (2) while $d$ is the number of parameters to fit, a ratio $n/d$ of 5 or 10 can yield good results. LIME by default sets $n$ to be in the thousands independently of $d$, which can be prohibitive for LLMs. Secondly, we limit the number of units $K$ that can be perturbed simultaneously to a small integer. This concentrates the smaller number of perturbations on inputs that are closer to $x^{o}$, which aligns with work showing that doing so improves the fidelity of attributions Tan et al. (2023). LIME by contrast samples the number of units to perturb uniformly from $\{1,\dots,d\}$ Mardaoui and Garreau (2021).
L-SHAP: This variation of SHAP by Chen et al. (2019) limits perturbations to a local neighborhood around a unit of interest, overcoming SHAP’s exponential complexity. We further extend L-SHAP by also limiting the number of simultaneously perturbed units (as in C-LIME). The equation for the SHAP score computation is given in (3). Please see Appendix B.3 for further details.
Multi-level explanations
The multi-level approach is essential for obtaining fine-grained attributions to the most important units in the context without a drastic increase in the computational cost. Once we have computed attribution scores at a given level, we refine the input units and repeat the process (feedback path at the bottom of Figure 2), generating multi-level explanations. For example, given sentence-level attributions, we obtain word-level attributions for a few sentences and keep attributions at the sentence level for the remainder to avoid introducing too many new units.
The few units to be refined are selected by Algorithm 1, [^3] Users can also manually select the units to be refined. which refines units with scores larger than a predefined threshold $\phi$ and that are among the $k$ units with the highest scores. Our multi-level approach decreases explanation cost by avoiding attribution of scores to less relevant finer units. See Appendix C.3 for the used hyperparameters.
Algorithm 1 Unit Refinement
1:$x_{1},\dots,x_{d}$triangleright Current units   $k\leq d$triangleright Maximum # units to refine   $\phi\in[-1,1]$triangleright Significance threshold
2:Compute attribution scores $\xi$ using (2) or (3).
3:$\psi\leftarrow 2\frac{\xi-\min_{s}\xi_{s}}{\max_{s}\xi_{s}-\min_{s}\xi_{s}}-1$triangleright Normalize scores
4:Compute Top-$k$($\psi$) $\triangleright$ $k$ largest $\psi_{s}$